The volume v ,of a liquid that flows through a pipe in time t,is given by
V/t =πpr^6/8cl

Where p,is pressure different between the ends of pipe of raidus r and length ,l is a constant
1.Determine the base units of c

2.Hence suggest what the quantity c represent
1
Expert's answer
2021-07-13T11:51:45-0400

Answer

By putting unit of each quantity in given formula

"\frac{m^3}{sec}=(\frac{N}{m^2}×m^6) /[c ]m"

[c] represents unit of "c"

So

"[c]=\frac{N×sec×m^6}{m^2×m^3×m}"


"[c]=\frac{N×sec×m^6}{m^6}"


"[c]=N-sec"


Its represent impulse.